OpenStudy (anonymous):

in calculus, what does the d/dx and d/dt mean? what's the difference? i am starting integration and am confused as to what the problem is asking for

OpenStudy (amistre64):

they voth mean basically the same theing, it is a notation that reads: the derivative with respect to "variable".

OpenStudy (nowhereman):

You could also say derivation in direction of some dimension.

OpenStudy (amistre64):

dy/dt tends to mean: the derivative of y with respect to t (for time)

OpenStudy (anonymous):

so what if it's dx/dt = 20/sqrt(t)? then it asks to find the antiderivative given x(1) = 40

OpenStudy (amistre64):

all it means is that the amount of change in the top part depends on the amount of change that occurs to the variable inthe bottom part

OpenStudy (amistre64):

initial conditions helpt to anchor the family of curves that is produced to a single curve.

OpenStudy (anonymous):

dy/d(something) just means the rate of change of 'y'(here) with respect to 'something'

OpenStudy (amistre64):

x' = 20/sqrt(t) means that there is a function X(t) that this came from.....and they want you to know where it came from...

OpenStudy (nowhereman):

Yes but as integration is the inverse operation of differentiation he is right, that in order to find x(t) he has to integrate dx/dt which in that example is 20/sqrt(t)

OpenStudy (amistre64):

take out the constant, and change the radical to an appropriate exponent

OpenStudy (anonymous):

so is dx/dt the derivative of a function called x? which in this case would be 20/sqrt(t)?

OpenStudy (nowhereman):

You should also know, that 1/sqrt(t) is the the derivative of 2\sqrt(t)

OpenStudy (amistre64):

20 times the integral of 1/t^(1/2)

OpenStudy (amistre64):

20 times the integral of t^(-1/2)

OpenStudy (amistre64):

20 (1/2)(sqrt(t) = 10sqrt(t)

OpenStudy (anonymous):

my book says the answer is 40*sqrt(t)

OpenStudy (amistre64):

poke: yes, dx/dt is the derivative of a funtion X(t) find X(t) and input your initial condition that x(1) = 40

OpenStudy (amistre64):

10sqrt(1) + C = 40

OpenStudy (amistre64):

if we derive 10sqrt(t) we get -> 10/2 sqrt(t)..right? so yeah, my mistake :)

OpenStudy (amistre64):

ahhh.... i see my mistake :) i did it in me head for starters. the reciprocal of 1/2 = 2..... 20(2) sqrt(t) = 40sqrt(t)

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

now this X(t) function is floating around....it needs an anchor...it needs a +C to hold it in place.

OpenStudy (amistre64):

40sqrt(t) + C is what you want.... plug in your initial condition that X(1) = 40 and solve for "C" then you have the right answer

OpenStudy (anonymous):

so when you pull the 20 out originally you have 20*(1/sqrt(t))... then the antideriv of that is 20*(2/sqrt(t))?

OpenStudy (amistre64):

poke: yes..

OpenStudy (anonymous):

so why isnt the final answer 40/sqrt(t)??

OpenStudy (amistre64):

you add one to the exponent, then put that number as the denominator and also use that number as the new exponent. -1/2 + 2/2 = 1/2

OpenStudy (amistre64):

40t^(1/2) = 40sqrt(t)

OpenStudy (amistre64):

t^(-1/2 + 2/2) 20t^(1/2) 20 * ---------------- = --------- = 40t^(1/2) -1/2 + 2/2 1/2

OpenStudy (amistre64):

which equals 40sqrt(t)

OpenStudy (anonymous):

ok... you changed the 1/sqrt(t) to 2/sqrt(t) and i got lost at that step. it makes sense now.

OpenStudy (amistre64):

its just reversing the steps of taking the derivative

OpenStudy (anonymous):

so basically when it says like dy/dt that means there is a function called y and the variable used will be t? such as dy/dt = 5t-4?

OpenStudy (amistre64):

thats correct

OpenStudy (amistre64):

but we have to finish our problem here..... tell me, do these curves have the same derivative: y= 2x+4 and y= 2x-17

OpenStudy (anonymous):

no

OpenStudy (amistre64):

take the derivative of both of them...what do you get?

OpenStudy (anonymous):

one is x^2 +4x and the other is x^2-17x

OpenStudy (amistre64):

not the integral....the derivative...

OpenStudy (anonymous):

o... both would be 2

OpenStudy (amistre64):

thats right.... then how do we know which curve is our answer when we suit them up with the integrals?

OpenStudy (amistre64):

dy/dx = 2 y = 2x + C right?

OpenStudy (amistre64):

where C can be any constant, any number

OpenStudy (anonymous):

right

OpenStudy (amistre64):

our integral is floating around..it needs to be anchored to a specific point in order to help us out

OpenStudy (anonymous):

in the example i originally gave the C would be 0. right?

OpenStudy (amistre64):

we have your problem.. X = 40sqrt(t) + C ....... how do we find the right equation of the curve? when we are given the X(1) = 40?

OpenStudy (anonymous):

u would do 40sqrt(1) + C = 40

OpenStudy (amistre64):

40 = 40sqrt(1) + C 40 = 40 + C 0 = C .... so yes, in this probelm C would be zero

OpenStudy (amistre64):

what would our answer be if we wanted X(1) = 20?

OpenStudy (anonymous):

1/2?

OpenStudy (amistre64):

20 = 40sqrt(1) + C 20 = 40 + C 20 -40 = C -20 = C X = 40sqrt(t) - 20 is the curve wed want right? you see how I did that?

OpenStudy (anonymous):

i mean -20

OpenStudy (anonymous):

ya. i divided it instead

OpenStudy (amistre64):

:) happens to me all the time lol

OpenStudy (amistre64):

does this make better sense to you now?

OpenStudy (anonymous):

ya. thanks for your help. i have to run off to class and take a quiz over this stuff. i should be fine now. thanks again!

OpenStudy (amistre64):

good luck, Ciao! :)